Let T : R2-+ R2be a linear operator with 

matrix [ (w.r.t. the standard basis).

[ 7 1

-1 -1]

Use Cayley-Hamilton theorem to 

check whether T is invertible or not. If T is 

invertible, obtain T-1(x, y) for 

(x,y) E R2 . If T is not invertible, obtain 

the minimum polynomial of T.


1
Expert's answer
2021-02-19T15:15:21-0500

Let the associated matrix be denoted by A. So we have that, "A = 
\begin{bmatrix}
 7 & 1 \\
 -1& -1
\end{bmatrix}"

"|A- \lambda I| = \begin{vmatrix}
 7- \lambda & 1 \\
 -1 & -1- \lambda
\end{vmatrix}
= (7- \lambda) (-1-\lambda)- 1(-1)"

= "\lambda^2 - 6\lambda - 6"

So we have that the characteristics polynomial is "\lambda^2 - 6\lambda - 6" . But Cayley Hamilton's theorem states that every square matrix satisfies it characteristics equation. So we check whether "A^2 - 6 A -6I = 0"

"A^2 =
\begin{bmatrix}
48 &6\\
-6&0
\end{bmatrix}\\
-6A =
\begin{bmatrix}
-42&-6\\
6&6
\end{bmatrix}\\
-6I=
\begin{bmatrix}
-6&0\\
0&-6
\end{bmatrix}"

So we have that "A^2 -6A -6I = 
\begin{bmatrix}
0&0\\
0&0
\end{bmatrix}"

as desired. So we have that A is invertible and hence T is invertible.

Since "A^2-6A-6I =0"

Multiplying through by "A^{-1}" , we have

"A-6I-6A^{-1}= 0\\
6A^{-1}= A-6I\\
A^{-1}= \frac{1}{6}(A-6I)"

"= \frac{1}{6}(\begin{bmatrix} 7&1\\-1&-1\end{bmatrix} - \begin{bmatrix} 6&0\\ 0&6 \end{bmatrix})"

"= \begin{bmatrix}
\frac{1}{6}&\frac{1}{6}\\ \frac{-1}{6}&\frac{-7}{6}
\end{bmatrix}"

Hence "T^{-1} =( \frac{1}{6}x+ \frac{1}{6}y, \frac{-1}{6}x -\frac{7}{6}y)"